OpenStudy (anonymous):

Which shows this equation in standard form? standard form please 6 – 2(x – y) = 3x + 5

OpenStudy (anonymous):

bootay

OpenStudy (anonymous):

are you solving for x? @labriya

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@labriya okay so we're solving for x 6 - 2 (x - y) = 3x + 5 add -3x to both sides -2x + 2 + 6 + -3x = 3x + 5 + -3x -5x + 2y + 6 = 5 add -6 to both siddes -5x + 2y + 6 + -6 = 5 + -6 -5x + 2y = -1 add -2y to both sides -5x + 2y + -2y = -1 + -2y -5x = -2y - 1 divide both sides by -5 -5x/-5 = -2y -1/ -5 answer -> x = 2/5y + 1/5

OpenStudy (gorv):

standard form is y=ax+b

OpenStudy (gorv):

u need to solve for y